When the interest rate is 10 percent, what is the price of a bond that matures in one year, has a $300 face value, and has a coupon of $20 per year?

$93.24

$266.11

$272.72

$290.90

$282.64
Answer #1

Total payment after 1 yr = 300 + 20 = 320

int rate = 10%

P = F/(1+i)^t

P = 320 / (1+0.1) = 290.90